OpenStudy (anonymous):

how does y=(x)/(x^2+1) have origin symmetry using the algebraic method ? I inputted (-x) and (-y) and it was different from the original equation.

OpenStudy (anonymous):

replace \(x\) by \(-x\) and you get \[\frac{-x}{(-x)^2+1}\] is a start

OpenStudy (anonymous):

then since \((-x)^2=x^2\) you have \[\frac{-x}{x^2+1}\] which is in fact \(-\frac{x}{x^2+1}=-y\)

OpenStudy (anonymous):

Omg. Thank you. So stupid of me. I really appreciate you answering my question !